Sucrose, C12H22O11, is the common table sugar. If 1.054 g of sucrose are added to a cup of tea how many molecules of sucrose will be in the cup?

A. 1.854 x 1021 molecules

B. 3.079 x 10-3 molecules

C. 1.251 x 1021 molecules

D. 2.078 x 10-3 molecules
Answer #1

1 mole of any substance is equal to 6.023 × 10^23 particles.

1 mole of sucrose = 6.023. × 10^23 molecules.

Number of moles = mass/molar mass

Molar mass of C12H22O11 = 342.297 g/mol

Mass of C12H22O11 = 1.054 g

Moles of C12H22O11 = 1.054/342.297

= 0.0031 moles

Number of molecules of C12H22O11 = 0.0031 × 6.023 × 10^23

= 1.867 × 10^21 molecules.

Answer -- A. 1.854 × 10^21 molecules